# The N-prime graph and the Subgroup Isomorphism Problem

### v1

@@ 4. The N-Prime Graph Question: ˜C
Let $p$ and $q$ be primes in $\pi(G)$ and, by means of contradiction, suppose that $q\to p$ is an arc of $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ but not of $\Gamma_{\rm{N}}(G)$. Then, taking into account Theorem˜3.5, $G$ is not isomorphic to the symmetric group ${\rm S}_{n}$, so we have $G\cong{\rm A}_{n}$ unless $n=6$ (and $|G|$ divides $4|{\rm A}_{n}|$). Moreover, $G$ does not have elements of order $pq$ and hence, by the main theorem of [3], neither does $V({\mathbb{Z}}G)$. Thus $V({\mathbb{Z}}G)$ has a subgroup $T=\left\langle u\right\rangle\rtimes\left\langle v\right\rangle$ with $|u|=p$, $|v|=q$ and $u^{v}=u^{r}$ for a suitable integer $r$ such that $r\not\equiv 1\,\,{\rm mod}\,\,\,p$ and $r^{q}\equiv 1\,\,{\rm mod}\,\,\,p$.
@@
By the Cohn-Livingstone Theorem, ${\rm A}_{n}$ has then an element $a$ of order $p$; in particular we have $n\geq p$, and we may take $a=(1,2,\dots,p)$. If $n\geq p+2$, then all the elements of ${\rm A}_{n}$ with the same cyclic type as $a$ lie in a single conjugacy class of ${\rm A}_{n}$, hence $a^{r}$ is conjugate to $a$ in ${\rm A}_{n}$. However, if $q$ is odd, then the same conclusion holds also for $n\in\{p,p+1\}$, because the map $x\mapsto x^{r}$ on ${\rm A}_{n}$ induces a permutation of odd order on the set $\mathcal{C}$ of conjugacy classes of elements having order $p$ in ${\rm A}_{n}$ and, in this case, we have $|\mathcal{C}|=2$. Now, if $a^{r}$ is conjugate to $a$ in ${\rm A}_{n}$, then we can find an element $b\in{\bf N}_{{\rm A}_{n}}(\langle a\rangle)$ with $|b|=q$, against our assumption. We conclude that $n$ lies in $\{p,p+1\}$ and $q=2$. Moreover, $T$ is a dihedral group of order $2p$ and $a$ is not conjugate to $a^{-1}$ in ${\rm A}_{n}$, i.e., $a$ is not a real element of ${\rm A}_{n}$. Note that the latter property does not hold if $p\equiv 1\,\,{\rm mod}\,\,\,4$, hence $n$ cannot be $6$ and we have $G\cong{\rm A}_{n}$.
@@
Since the element $a\in G$ is not real, we can take an irreducible character $\chi$ of $G$ such that $\chi(a)\not\in\mathbb{R}$. Still denoting by $\chi$ the linear expansion of this character to ${\mathbb{Z}}G$, we have that the restriction $\chi_{T}$ is a character of $T$.
Fix a complex primitive $p$-th root of the unity $\zeta$ and consider the character $\psi$ of $T$ induced by the linear character of $\left\langle u\right\rangle$ mapping $u$ to $\zeta$. Then $\psi(u^{i})=\zeta^{i}+\zeta^{-i}$ and $\psi(u^{i}v)=0$ for every $i\in\{0,1,\dots,p-1\}$. As $\chi_{T}$ and $\psi$ are characters of $T$, the inner product $\left\langle\psi,\chi_{T}\right\rangle$ is a non-negative integer, so its imaginary part is 0.
@@
But $G$ has exactly two conjugacy classes of elements of order $p$, represented by $a$ and $a^{-1}$. By the Berman-Higman Theorem and [12, Theorem 2.3], if $1\leq i\leq p-1$ and $g\in G$ is such that $\varepsilon_{g}(u^{i})\neq 0$, then $g$ is conjugate to $a$ or $a^{-1}$. Therefore $\varepsilon_{a}(u^{i})+\varepsilon_{a^{-1}}(u^{i})=1$ and
$\chi(u^{i})=\varepsilon_{a}(u^{i})\chi(a)+\varepsilon_{a^{-1}}(u^{i})\overline{\chi(a)}=\varepsilon_{a}(u^{i})\chi(a)+(1-\varepsilon_{a}(u^{i}))\overline{\chi(a)}.$
Thus, if $x$ is the imaginary part of $\chi(a)$, then $x\neq 0$ and the imaginary part of $\left\langle\psi,\chi_{T}\right\rangle$ is
$0=\frac{1}{2p}\sum_{i=1}^{\frac{p-1}{2}}(\zeta^{i}+\zeta^{-i})(\varepsilon_{a}(u^{i})x-(1-\varepsilon_{a}(u^{i}))x)=\frac{x}{2p}\sum_{i=1}^{\frac{p-1}{2}}(\zeta^{i}+\zeta^{-i})(2\varepsilon_{a}(u^{i})-1).$
Since $\{\zeta^{i}+\zeta^{-i}:1\leq i\leq\frac{p-1}{2}\}$ is a basis of ${\mathbb{Q}}(\zeta+\zeta^{-1})$ over $\mathbb{Q}$} and each $u^{i}\in{\mathbb{Z}}G$, it follows that $\varepsilon_{a}(u^{i})\in{\mathbb{Z}}$ and $2\varepsilon_{a}(u^{i})=1$, a contradiction which completes the proof.
@@
    The set $\{\chi(a)\mid a\in A\}$ consists of complex numbers that are linearly independent over $\mathbb{Q}$}.
@@
Every character $\varphi\in{\rm Irr}(G)$ extends uniquely to a linear map ${\mathbb{C}}G\to{\mathbb{C}}$, which we also denote by $\varphi$. As $u$ and $u^{r}$ are conjugate in $V({\mathbb{Z}}G)$, and hence in ${\mathbb{C}}G$, it follows that $\varphi(u)=\varphi(u^{r})$. Also, it is not difficult to see that we have $\varphi(u^{r})=\sigma(\varphi(u))$. Similarly, $\chi$ extends uniquely to a linear map ${\mathbb{C}}G\to{\mathbb{C}}$, which we also denote $\chi$. Using the expression of $\chi$ as a linear combination of irreducible characters of $G$ we obtain $\sigma(\chi(u))=\chi(u^{r})=\chi(u)$. On the other hand, by the Berman-Higman Theorem and [12, Theorem 2.3], $\varepsilon_{g}(u)=0$ for every $g\in G$ of order different from $p$. Hence, since $\chi$ is a class function, we have
@@
Since ${\rm{PSL}}_{2}(2^{2})$ is isomorphic to ${\rm A}_{5}$, which has been already treated in Theorem˜4.1, we can assume that $r$ is an odd prime. We start by recalling that the order of $S$ for $r\neq 2$ is ${r^{f}(r^{f}-1)(r^{f}+1)}/2$, and the order of ${\rm{Out}}(S)$ is $2f$; in fact, the automorphism group of $S$ is isomorphic to $S\left\langle\delta,\phi\right\rangle$, where $\delta$ is a diagonal automorphism such that $\delta^{2}$ is an inner automorphism, and $\phi$ is a field automorphism of order $f$ (see for instance [28]). Therefore $\pi(G)=\pi(S)=\pi^{-}\cup\pi^{+}\cup\{r\}\cup\{2\}$, where $\pi^{-}$ and $\pi^{+}$ are the sets of odd prime divisors of $r^{f}-1$ and of $r^{f}+1$, respectively.
By [17, II.8.27], the group $S$ has dihedral subgroups both of order $r^{f}-1$ and of order $r^{f}+1$; also, it has subgroups that are Frobenius groups with an elementary abelian kernel of order $r^{f}$ and cyclic complements of order $(r^{f}-1)/2$. In view of this fact, if $r^{f}\equiv 1\,\,{\rm mod}\,\,\,4$, then $\Gamma_{\rm{N}}(S)$ contains the following arcs: $s\leftrightarrows t$ for any choice of $s$ and $t$ in $\{2\}\cup\pi^{-}$ or in $\pi^{+}$, and $2\rightarrow s$ for any $s\in\{r\}\cup\pi^{+}$. Similarly, if $r^{f}\equiv-1\,\,{\rm mod}\,\,\,4$, then $\Gamma_{\rm{N}}(S)$ contains the arcs $s\leftrightarrows t$ for any choice of $s$ and $t$ in $\{2\}\cup\pi^{+}$ or in $\pi^{-}$, and $2\rightarrow s$ for any $s\in\pi^{-}$. Moreover, in both cases $\Gamma_{\rm{N}}(S)$ has the arcs $s\rightarrow r$ where $s\in\pi^{-}$ is a divisor of r1r-1.
Now, all these arcs are clearly in $\Gamma_{\rm{N}}(G)$, and also $2\rightarrow r$ is in $\Gamma_{\rm{N}}(G)$ unless $r^{f}\equiv-1\,\,{\rm mod}\,\,\,4$ and $G=S$; in fact, it is not difficult to check that $G$ has always a (possibly abelian) subgroup of the form $C_{r}\rtimes C_{2}$ whenever $G$ is not simple. Furthermore, every double arc in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ is also in $\Gamma_{\rm{N}}(G)$ by [24, Theorem 5.5] and Remark˜3.1. Since, by obvious arithmetical reasons, $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ cannot contain arcs of the kind $s\rightarrow r$ for any $s\in\pi^{+}$, or $s$ in $\pi^{-}$ not dividing r1r-1, the desired conclusion will be obtained by proving the following properties.
@@
    $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ does not contain any arc of the form $q\rightarrow p$, for $p\in\pi^{+}$ and $q\in\pi^{-}$ or vice-versa.
@@
So, let us define $\xi$ to be either $\theta_{1}$ or $\chi_{1}$, depending on whether we are in the case $p\in\pi^{+}$ and $q\in\pi^{-}$ or vice-versa, and assume that $\xi$ does not extend to $G$. Let $\psi$ be an irreducible character of $G$ whose restriction to $S={\rm{PSL}}_{2}(r^{2})$ has $\xi$ as a constituent. Since the diagonal automorphism $\delta$ lies in the inertia subgroup of $\xi$ in ${\rm Aut}(S)$ (see [28, Page 2]), we have that the restriction of $\psi$ to $S$ splits as $\xi+\xi^{\phi}$. Now, the automorphism $\phi$ induces a permutation on the set $\mathcal{C}=\{C_{1},\ldots,C_{t}\}$ of $S$-conjugacy classes of elements of order $p$, where the class $x^{S}\in\mathcal{C}$ is mapped to $(x^{r})^{S}$, and we consider a set $\mathcal{F}=\{C_{1},\ldots,C_{s}\}$ of representatives for the orbits of this action. Note that, since $\delta$ fixes all the elements of $\mathcal{C}$, a set $\{x_{1},\ldots,x_{s}\}$ of representatives for the $S$-classes in $\mathcal{F}$ is also a set of representatives for the $G$-conjugacy classes of elements of order $p$.
We claim that the complex numbers in the set $\{\psi(x_{1}),\ldots,\psi(x_{s})\}$ are linearly independent over $\mathbb{Q}$}. In fact, assume that we have
@@
and this is a linear combination with rational coefficients of values of $\xi$ on a set of representatives for the $S$-conjugacy classes in $\mathcal{C}$ (indeed, if $p$ lies in $\pi^{-}$ then we have $\xi(x_{j})=\xi(x_{j}^{r})$ for every $j\in\{1,\ldots s\}$, whereas if $p$ lies in $\pi^{+}$ then $\xi(x_{j})\neq\xi(x_{j}^{r})$ for every $j\in\{1,\ldots s\}$). Since we know this values are linearly independent over $\mathbb{Q}$}, it easily follows that $q_{j}=0$ for all $j\in\{1,\ldots,s\}$. Moreover, for every $j\in\{1,\dots,s\}$, we have that $q$ does not divide $|{\mathbb{Q}}_{p}:{\mathbb{Q}}(\psi(x_{j})+\psi(x_{j}^{r}))|$ (which is either $2$ or $4$, depending on whether $p$ is in $\pi^{-}$ or in $\pi^{+}$, respectively), and again we are done by Lemma 4.2. This concludes the proof of the claim in (a).
@@
5. Metacyclic subgroups of solvable groups: ˜D and ˜E
We work next toward an improvement of ˜B in the spirit of the Subgroup Isomorphism Problem, as described in the introduction. We will first prove that if $G$ is a finite solvable group and $V({\mathbb{Z}}G)$ has a subgroup isomorphic to a Frobenius group of the form $C_{p}\rtimes C_{q^{k}}$, then so does $G$ (˜D of the introduction).
@@ 5. Metacyclic subgroups of solvable groups: ˜D and ˜E
Let $H$ be a minimal counterexample to the statement, and let $P=\left\langle a\right\rangle$ be a Sylow $p$-subgroup of $H$. Observe that, by the Frattini argument, we have $H=N{\bf N}_{H}(P)$ and, by [9, Theorem 5.2.3], $N=[N,P]\times{\bf C}_{N}(P)$; therefore, setting $M=[N,P]$ and $C={\bf C}_{N}(P)$, we have $H=MC{\bf N}_{H}(P)=M{\bf N}_{H}(P)$. In particular, $M$ and $MP$ are normal subgroups of $H$, because they are normalized by both $M$ and ${\bf N}_{H}(P)$. Since $M\cap{\bf N}_{H}(P)={\bf N}_{M}(P)={\bf C}_{M}(P)=M\cap C=1$, we see that ${\bf N}_{H}(P)$ is in fact a complement for $M$ in $H$.
@@
As the next step, we will count the elements of order $q^{k}$ in $H$ that normalize some Sylow $p$-subgroup of $H$. Note that, if $y$ is such an element, then $y^{q^{k-1}}$ normalizes a Sylow $p$-subgroup of $H$ and hence it does not lie in $M$; it follows that $\left\langle y\right\rangle\cap M=1$. Observe that every Sylow $p$-subgroup of $H$ is contained in $MP$ and, again because $MP$ is a Frobenius group with complement $P$, we get $|{\rm Syl}_{p}(H)|=|M|$. In fact, we have ${\rm Syl}_{p}(H)=\{P^{m}\mid m\in M\}$. It follows that the normalizer of every Sylow $p$-subgroup of $H$ is isomorphic to the Frobenius group $C_{p}\rtimes C_{q^{k}}$, and therefore the number of subgroups of order $q^{k}$ of $H$ which normalize a given Sylow $p$-subgroup of $H$ is exactly $p$. We also observe that, by Theorem 15.16 of [18], for every element $y\in H\setminus M$ having order $q^{k}$ we get $|M|=|{\bf C}_{M}(y)|^{q^{k}}$; in particular, $n=|{\bf C}_{M}(y)|$ does not depend on the choice of $y$ among the elements of order $q^{k}$ in $H\setminus M$. Now, we claim that each of these elements of order $q^{k}$ normalizes precisely $n$ distinct Sylow $p$-subgroups of $H$. In fact, assume that $y\in H$ of order $q^{k}$ normalizes $P_{0}\in{\rm Syl}_{p}(H)$. If, for $m\in M$, we have $y\in{\bf N}_{H}(P_{0}^{m})$, then $m^{-1},y^{-1}]$ lies in ${\bf N}_{H}(P_{0})y^{-1}\cap M={\bf N}_{H}(P_{0})\cap M=1$ and so $m$ lies in ${\bf C}_{M}(y)$. On the other hand, it is clear that if $y$ centralizes $m$ then $y$ normalizes $P_{0}^{m}$, and our claim follows.
@@
We are ready to conclude the count mentioned in the previous paragraph. Denoting by $\mathcal{C}$ the set of subgroups of order $q^{k}$ in $H$ that normalize some Sylow $p$-subgroup of $H$, consider the bipartite graph whose vertex set is ${\rm Syl}_{p}(H)\cup\mathcal{C}$ (with ${\rm Syl}_{p}(H)$ and $\mathcal{C}$ being the two parts of the graph), and where the vertices $P_{0}\in{\rm Syl}_{p}(H)$ and $Q_{0}\in\mathcal{C}$ are adjacent if and only if $Q_{0}$ normalizes $P_{0}$. By the discussion above, the number of edges in this graph can be computed both as $|M|p$ (which is the number of vertices in the part ${\rm Syl}_{p}(H)$ multiplied by the number of neighbors of each vertex in ${\rm Syl}_{p}(H)$) and $|\mathcal{C}|n$ (obtained similarly, from the point of view of the part $\mathcal{C}$). So, we have $|\mathcal{C}|=|M|p/n=n^{q^{k}-1}p$. As each subgroup of order $q^{k}$ contains $q^{k-1}(q-1)$ elements of order $q^{k}$, we conclude that the number of elements of order $q^{k}$ in $H$ which normalize some Sylow $p$-subgroup of $H$ is given by
$|\mathcal{C}|q^{k-1}(q-1)=n^{q^{k}-1}pq^{k-1}(q-1).$

### v2

@@ 4. The N-Prime Graph Question: Theorem C
Let $p$ and $q$ be primes in $\pi(G)$ and, by means of contradiction, suppose that $q\to p$ is an arc of $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ but not of $\Gamma_{\rm{N}}(G)$. Then, taking into account Theorem 3.5, $G$ is not isomorphic to the symmetric group ${\rm S}_{n}$, so we have $G\cong{\rm A}_{n}$ unless $n=6$ (and $|G|$ divides $4|{\rm A}_{n}|$). Moreover, $G$ does not have elements of order $pq$ and hence, by the main theorem of [3], neither does $V({\mathbb{Z}}G)$. Thus $V({\mathbb{Z}}G)$ has a subgroup $T=\left\langle u\right\rangle\rtimes\left\langle v\right\rangle$ with $|u|=p$, $|v|=q$ and $u^{v}=u^{r}$ for a suitable integer $r$ such that $r\not\equiv 1\,\,{\rm mod}\,\,\,p$ and $r^{q}\equiv 1\,\,{\rm mod}\,\,\,p$.
@@
By Corollary 4.1 of [6], ${\rm A}_{n}$ has then an element $a$ of order $p$; in particular we have $n\geq p$, and we may take $a=(1,2,\dots,p)$. If $n\geq p+2$, then all the elements of ${\rm A}_{n}$ with the same cyclic type as $a$ lie in a single conjugacy class of ${\rm A}_{n}$, hence $a^{r}$ is conjugate to $a$ in ${\rm A}_{n}$. However, if $q$ is odd, then the same conclusion holds also for $n\in\{p,p+1\}$, because the map $x\mapsto x^{r}$ on ${\rm A}_{n}$ induces a permutation of odd order on the set $\mathcal{C}$ of conjugacy classes of elements having order $p$ in ${\rm A}_{n}$ and, in this case, we have $|\mathcal{C}|=2$. Now, if $a^{r}$ is conjugate to $a$ in ${\rm A}_{n}$, then we can find an element $b\in{\bf N}_{{\rm A}_{n}}(\langle a\rangle)$ with $|b|=q$, against our assumption. We conclude that $n$ lies in $\{p,p+1\}$ and $q=2$. Moreover, $T$ is a dihedral group of order $2p$ and $a$ is not conjugate to $a^{-1}$ in ${\rm A}_{n}$, i.e., $a$ is not a real element of ${\rm A}_{n}$. Note that the latter property does not hold if $p\equiv 1\,\,{\rm mod}\,\,\,4$, hence $n$ cannot be $6$ and we have $G\cong{\rm A}_{n}$.
@@
Since the element $a\in G$ is not real, we can take an irreducible character $\chi$ of $G$ such that $\chi(a)\not\in\mathbb{R}$, and we denote by $x$ the (non-zero) imaginary part of $\chi(a)$. On the other hand, as $u$ is conjugate to $u^{-1}$ in $T$, the linear expansion $\widehat{\chi}$ of $\chi$ to ${\mathbb{Z}}G$ takes a real value on $u$.
@@
Now, $G$ has exactly two conjugacy classes of elements of order $p$, represented by $a$ and $a^{-1}$. By the Berman-Higman Theorem and [13, Theorem 2.3], if $g\in G$ is such that $\varepsilon_{g}(u)\neq 0$, then $g$ is conjugate to $a$ or $a^{-1}$. Therefore $\varepsilon_{a}(u)+\varepsilon_{a^{-1}}(u)=1$ and
${\widehat{\chi}}(u)=\varepsilon_{a}(u)\chi(a)+\varepsilon_{a^{-1}}(u)\overline{\chi(a)}=\varepsilon_{a}(u)\chi(a)+(1-\varepsilon_{a}(u))\overline{\chi(a)}=\varepsilon_{a}(u)(\chi(a)-{\overline{\chi(a)}})+{\overline{\chi(a)}}.$
Comparing the imaginary parts of the two sides, we get $0=x\cdot(2\varepsilon_{a}(u)-1)$; but $x\neq 0$, and $\varepsilon_{a}(u)$ is then an integer such that $2\varepsilon_{a}(u)-1=0$. This is the final contradiction which completes the proof.
@@
    The complex numbers $\chi(a)$, for $a\in A$, are linearly independent over $\mathbb{Q}$}.
@@
Every character $\varphi\in{\rm Irr}(G)$ extends uniquely to a linear map ${\mathbb{C}}G\to{\mathbb{C}}$, which we also denote by $\varphi$. As $u$ and $u^{r}$ are conjugate in $V({\mathbb{Z}}G)$, and hence in ${\mathbb{C}}G$, it follows that $\varphi(u)=\varphi(u^{r})$. Also, it is not difficult to see that we have $\varphi(u^{r})=\sigma(\varphi(u))$. Similarly, $\chi$ extends uniquely to a linear map ${\mathbb{C}}G\to{\mathbb{C}}$, which we also denote $\chi$. Using the expression of $\chi$ as a linear combination of irreducible characters of $G$ we obtain $\sigma(\chi(u))=\chi(u^{r})=\chi(u)$. On the other hand, by the Berman-Higman Theorem and [13, Theorem 2.3], $\varepsilon_{g}(u)=0$ for every $g\in G$ of order different from $p$. Hence, since $\chi$ is a class function, we have
@@
Since ${\rm{PSL}}_{2}(2^{2})$ is isomorphic to ${\rm A}_{5}$, which has been already treated in Theorem 4.1, we can assume that $r$ is an odd prime. We start by recalling that the order of $S$ for $r\neq 2$ is ${r^{f}(r^{f}-1)(r^{f}+1)}/2$, and the order of ${\rm{Out}}(S)$ is $2f$; in fact, the automorphism group of $S$ is isomorphic to $S\left\langle\delta,\phi\right\rangle$, where $\delta$ is a diagonal automorphism such that $\delta^{2}$ is an inner automorphism, and $\phi$ is a field automorphism of order $f$ (see for instance [30]). Therefore $\pi(G)=\pi(S)=\pi^{-}\cup\pi^{+}\cup\{r\}\cup\{2\}$, where $\pi^{-}$ and $\pi^{+}$ are the sets of odd prime divisors of $r^{f}-1$ and of $r^{f}+1$, respectively.
By [18, II.8.27], the group $S$ has dihedral subgroups both of order $r^{f}-1$ and of order $r^{f}+1$; also, it has subgroups that are Frobenius groups with an elementary abelian kernel of order $r^{f}$ and cyclic complements of order $(r^{f}-1)/2$. In view of this fact, if $r^{f}\equiv 1\,\,{\rm mod}\,\,\,4$, then $\Gamma_{\rm{N}}(S)$ contains the following arcs: $s\leftrightarrows t$ for any choice of $s$ and $t$ in $\{2\}\cup\pi^{-}$ or in $\pi^{+}$, and $2\rightarrow s$ for any $s\in\{r\}\cup\pi^{+}$. Similarly, if $r^{f}\equiv-1\,\,{\rm mod}\,\,\,4$, then $\Gamma_{\rm{N}}(S)$ contains the arcs $s\leftrightarrows t$ for any choice of $s$ and $t$ in $\{2\}\cup\pi^{+}$ or in $\pi^{-}$, and $2\rightarrow s$ for any $s\in\pi^{-}$. Moreover, in both cases $\Gamma_{\rm{N}}(S)$ has the arcs $s\rightarrow r$ where $s\in\pi^{-}$ is a divisor of r1r-1.
Now, all these arcs are clearly in $\Gamma_{\rm{N}}(G)$, and also $2\rightarrow r$ is in $\Gamma_{\rm{N}}(G)$ unless $r^{f}\equiv-1\,\,{\rm mod}\,\,\,4$ and $G=S$; in fact, it is not difficult to check that $G$ has always a (possibly abelian) subgroup of the form $C_{r}\rtimes C_{2}$ whenever $G$ is not simple. Furthermore, every double arc in $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ is also in $\Gamma_{\rm{N}}(G)$ by [26, Theorem 5.5] and Remark 3.1. Since, by obvious arithmetical reasons, $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ cannot contain arcs of the kind $s\rightarrow r$ for any $s\in\pi^{+}$, or $s$ in $\pi^{-}$ not dividing r1r-1, the desired conclusion will be obtained by proving the following properties.
@@
    $\Gamma_{\rm{N}}(V({\mathbb{Z}}G))$ does not contain any arc of the form $q\rightarrow p$, for any choice of a pair $(p,q)$ in $\pi^{+}\times\pi^{-}$ or in $\pi^{-}\times\pi^{+}$.
@@
So, let us define $\xi$ to be either $\theta_{1}$ or $\chi_{1}$, depending on whether we are in the case $p\in\pi^{+}$ and $q\in\pi^{-}$ or vice-versa, and assume that $\xi$ does not extend to $G$. Let $\psi$ be an irreducible character of $G$ whose restriction to $S={\rm{PSL}}_{2}(r^{2})$ has $\xi$ as a constituent. Since the diagonal automorphism $\delta$ lies in the inertia subgroup of $\xi$ in ${\rm Aut}(S)$ (see [30, Page 2]), we have that the restriction of $\psi$ to $S$ splits as $\xi+\xi^{\phi}$. Now, the automorphism $\phi$ induces a permutation on the set $\mathcal{C}=\{C_{1},\ldots,C_{t}\}$ of $S$-conjugacy classes of elements of order $p$, where the class $x^{S}\in\mathcal{C}$ is mapped to $(x^{r})^{S}$, and we consider a set $\mathcal{F}=\{C_{1},\ldots,C_{s}\}$ of representatives for the orbits of this action. Note that, since $\delta$ fixes all the elements of $\mathcal{C}$, a set $\{x_{1},\ldots,x_{s}\}$ of representatives for the $S$-classes in $\mathcal{F}$ is also a set of representatives for the $G$-conjugacy classes of elements of order $p$.
We claim that the complex numbers $\psi(x_{1}),\ldots,\psi(x_{s})$ are linearly independent over $\mathbb{Q}$}. In fact, assume that we have
@@
and this is a linear combination with rational coefficients of values of $\xi$ on a set of representatives for the $S$-conjugacy classes in $\mathcal{C}$ (indeed, if $p$ lies in $\pi^{-}$ then we have $\xi(x_{j})=\xi(x_{j}^{r})$ for every $j\in\{1,\ldots s\}$, whereas if $p$ lies in $\pi^{+}$ then $\xi(x_{j})\neq\xi(x_{j}^{r})$ for every $j\in\{1,\ldots s\}$). Since we know these values are linearly independent over $\mathbb{Q}$}, it easily follows that $q_{j}=0$ for all $j\in\{1,\ldots,s\}$. Moreover, for every $j\in\{1,\dots,s\}$, we have that $q$ does not divide $|{\mathbb{Q}}_{p}:{\mathbb{Q}}(\psi(x_{j})+\psi(x_{j}^{r}))|$ (which is either $2$ or $4$, depending on whether $p$ is in $\pi^{-}$ or in $\pi^{+}$, respectively), and again we are done by Lemma 4.2. This concludes the proof of the claim in (a).
@@
5. Metacyclic subgroups of solvable groups: Theorem D and Theorem E
We work next toward an improvement of Corollary B in the spirit of the Subgroup Isomorphism Problem, as described in the introduction. We will first prove that if $G$ is a finite solvable group and $V({\mathbb{Z}}G)$ has a subgroup isomorphic to a Frobenius group of the form $C_{p}\rtimes C_{q^{k}}$, then so does $G$ (Theorem D of the introduction).
@@ 5. Metacyclic subgroups of solvable groups: Theorem D and Theorem E
Let $H$ be a minimal counterexample to the statement, and let $P=\left\langle a\right\rangle$ be a Sylow $p$-subgroup of $H$. Observe that, by the Frattini argument, we have $H=N{\bf N}_{H}(P)$ and, by [10, Theorem 5.2.3], $N=[N,P]\times{\bf C}_{N}(P)$; therefore, setting $M=[N,P]$ and $C={\bf C}_{N}(P)$, we have $H=MC{\bf N}_{H}(P)=M{\bf N}_{H}(P)$. In particular, $M$ and $MP$ are normal subgroups of $H$, because they are normalized by both $M$ and ${\bf N}_{H}(P)$. Since $M\cap{\bf N}_{H}(P)={\bf N}_{M}(P)={\bf C}_{M}(P)=M\cap C=1$, we see that ${\bf N}_{H}(P)$ is in fact a complement for $M$ in $H$.
@@
As the next step, we will count the elements of order $q^{k}$ in $H$ that normalize some Sylow $p$-subgroup of $H$. Note that, if $y$ is such an element, then $y^{q^{k-1}}$ normalizes a Sylow $p$-subgroup of $H$. Since, by the first paragraph of this proof, the normalizer in $M$ of any Sylow $p$-subgroup of $H$ is trivial, it follows that $\left\langle y\right\rangle\cap M=1$. Observe that every Sylow $p$-subgroup of $H$ is contained in $MP$ and, again because $MP$ is a Frobenius group with complement $P$, we get $|{\rm Syl}_{p}(H)|=|M|$. In fact, we have ${\rm Syl}_{p}(H)=\{P^{m}\mid m\in M\}$. It follows that the normalizer of every Sylow $p$-subgroup of $H$ is isomorphic to the Frobenius group $C_{p}\rtimes C_{q^{k}}$, and therefore the number of subgroups of order $q^{k}$ of $H$ which normalize a given Sylow $p$-subgroup of $H$ is exactly $p$. We also observe that, by Theorem 15.16 of [19], for every element $y\in H\setminus M$ having order $q^{k}$ we get $|M|=|{\bf C}_{M}(y)|^{q^{k}}$; in particular, $\ell=|{\bf C}_{M}(y)|$ does not depend on the choice of $y$ among the elements of order $q^{k}$ in $H\setminus M$. Now, we claim that each of these elements of order $q^{k}$ normalizes precisely $\ell$ distinct Sylow $p$-subgroups of $H$. In fact, assume that $y\in H$ of order $q^{k}$ normalizes $P_{0}\in{\rm Syl}_{p}(H)$. If, for $m\in M$, we have $y\in{\bf N}_{H}(P_{0}^{m})$, then $m^{-1},y^{-1}]$ lies in ${\bf N}_{H}(P_{0})y^{-1}\cap M={\bf N}_{H}(P_{0})\cap M=1$ and so $m$ lies in ${\bf C}_{M}(y)$. On the other hand, it is clear that if $y$ centralizes $m$ then $y$ normalizes $P_{0}^{m}$, and our claim follows.
@@
We are ready to conclude the count mentioned in the previous paragraph. Denoting by $\mathcal{C}$ the set of subgroups of order $q^{k}$ in $H$ that normalize some Sylow $p$-subgroup of $H$, consider the bipartite graph whose vertex set is ${\rm Syl}_{p}(H)\cup\mathcal{C}$ (with ${\rm Syl}_{p}(H)$ and $\mathcal{C}$ being the two parts of the graph), and where the vertices $P_{0}\in{\rm Syl}_{p}(H)$ and $Q_{0}\in\mathcal{C}$ are adjacent if and only if $Q_{0}$ normalizes $P_{0}$. By the discussion above, the number of edges in this graph can be computed both as $|M|p$ (which is the number of vertices in the part ${\rm Syl}_{p}(H)$ multiplied by the number of neighbors of each vertex in ${\rm Syl}_{p}(H)$) and $|\mathcal{C}|\ell$ (obtained similarly, from the point of view of the part $\mathcal{C}$). So, we have $|\mathcal{C}|=|M|p/\ell=\ell^{q^{k}-1}p$. As each subgroup of order $q^{k}$ contains $q^{k-1}(q-1)$ elements of order $q^{k}$, we conclude that the number of elements of order $q^{k}$ in $H$ which normalize some Sylow $p$-subgroup of $H$ is given by
$|\mathcal{C}|q^{k-1}(q-1)=\ell^{q^{k}-1}pq^{k-1}(q-1).$